OpenStudy (anonymous):

Is this correct? The quotient of f(x)= x-3 and g(x) = x-5 is 3/5 and the domain is [5, infinity)?

OpenStudy (anonymous):

The quotient should just be (x-3)/(x-5), since if we sub in x =1, we get (1-3)/(1-5) which equals -2/-4 and that is 1/2. So it wouldn't be 3/5. As for the domain for either of the functions, x should be all real values.

OpenStudy (anonymous):

it will be 2.. for domain\[-\infty \to +\infty\] if f(x) is divided by g(x)

OpenStudy (anonymous):

x can't equal 5

OpenStudy (anonymous):

x can equal 5, it would just mean that f(5) = 0.

OpenStudy (anonymous):

If x = 5 you would have 2/0 which is not possible

OpenStudy (anonymous):

because zero can't be in the denominator

OpenStudy (anonymous):

Wait, aren't they two separate functions?

OpenStudy (anonymous):

Oh my mistake, she was looking for the function of the quotient, not f(x) or g(x). Yeah, x cannot be 5

OpenStudy (anonymous):

so for interval notation is it \[[-5,\infty)\]

OpenStudy (anonymous):

That's only part of the solution

OpenStudy (anonymous):

actually...

OpenStudy (anonymous):

i mean \[[5,\infty)\]

OpenStudy (anonymous):

(-infinity, 5), (5, +infinity)

OpenStudy (anonymous):

That's the solution

OpenStudy (anonymous):

the quotient is equal to 1 the remainder is equal to 2 (x - 3)/(x - 5) = 1 + 2/(x-5)

OpenStudy (anonymous):

so is it a union for the domain?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry, i didn't mean to post that so many times.

OpenStudy (anonymous):

and this is all in parenthesis?

OpenStudy (anonymous):

yes, correct because 5 is not included in the solution.

OpenStudy (anonymous):

how do you know that the domain is a union and not one single set of intervals?

OpenStudy (anonymous):

Because the solution is incomplete without the other set. Providing only one set is incomplete.

OpenStudy (anonymous):

When you are providing a solution, you have to provide the complete solution.

OpenStudy (anonymous):

ok, i think i got it. it's basically saying that the domain is negative infinty, not including 5, and 5 to infinity

OpenStudy (anonymous):

Precisely

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

R\{5} is probably the most succinct way to write the domain using set notation

OpenStudy (anonymous):

okay, my teacher hasn't taught that yet, i'm not sure if she will, so i'll just use a union.